OpenStudy (anonymous):

Two thirds of all students at school are girls. Two thirds of all students have dark hair. If half the boys have dark hair, then what portion of the girls have black hair? Express your answer as a common fraction Two thirds of all students at school are girls. Two thirds of all students have dark hair. If half the boys have dark hair, then what portion of the girls have black hair? Express your answer as a common fraction @Mathematics

OpenStudy (anonymous):

1/2

OpenStudy (asnaseer):

call total number of students T. the girls = 2T/3, boys=T/3 dark haired = 2T/3 half boys have dark hair, so (1/2)*(number of boys)=(1/2)*(T/3)=T/6 but total students with dark hair is 2T/3, so remaining dark haired students (all girls) must be = (2T/3)-(T/6)=3T/6=T/2 total girls is 2T/3, so portion of girls that have dark hair = (1/2) / (2/3) = (1/2) * (3/2) = 3/4 so 3/4 of girls have dark hair

OpenStudy (anonymous):

what portion of the girls have black hair? that was the question what portion of girls have dark hair i said 1/2 portion of girls have dark hair

OpenStudy (anonymous):

sheg has it. imagine 6 students, 3 with dark hair, consisting of 4 girls, 2 boys. 1 boy has dark hair and therefore 2 girls (half the number of girls) have dark hair.

OpenStudy (anonymous):

so the answer is 1/2

OpenStudy (anonymous):

yes and hint for future questions: since the total was not given in the problem, it makes no difference what the total it. therefore feel free to make up simple numbers to work with. whatever answer you get ( if you do it right) will be correct.

OpenStudy (asnaseer):

satellite73 - if you use 6 as the total number of students then number with dark hair should be 4 (2/3 of 6)

OpenStudy (anonymous):

oh so i did. i thought it said "half the students have dark hair"

OpenStudy (anonymous):

satelite asnaseer is right we have taken total remaining students with dark hairs here

OpenStudy (anonymous):

you are correct and i am wrong. should be 4 with dark hair, therefore 1 by and 3 girls.

OpenStudy (anonymous):

*boy

OpenStudy (anonymous):

can someone answer my domino question i cant figure that out

OpenStudy (anonymous):

therefore 3 out of 4 girls have dark hair, and the answer is 3/4